OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = -x^4 + 4x^3 - 2x^2 + x + 1, from x = 0 to x =1 Options 1)0 2)1 3)2 4)7

OpenStudy (amistre64):

any ideas?

OpenStudy (amistre64):

an average rate of change is simply the slope of the line between 2 points ...

OpenStudy (anonymous):

i have no idea this problem makes no sense to me @amistre64

OpenStudy (amistre64):

so, you have no idea how to find the slope between 2 points?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

well, there is something called a slope formula. it is the change in y, divided by the change in x. let x move from a to b, then y moves from f(a) to f(b) and we get: \[slope=\frac{f(a)-f(b)}{a-b}\]

OpenStudy (anonymous):

how do i know what numbers to plug in?

OpenStudy (amistre64):

they define f(x) for you already, and tell you that x is from 0 to 1.

OpenStudy (amistre64):

\[\frac{f(0)-f(1)}{0-1}\]

OpenStudy (anonymous):

wouldnt that be -1

OpenStudy (amistre64):

dunno. i havent tried to work it yet :)

OpenStudy (amistre64):

f(0) = 1 f(1) = -1 + 4 - 2 + 1 + 1 = 3 1-3 --- = 3 0-1

OpenStudy (amistre64):

im going with no, its not -1

OpenStudy (anonymous):

so its 3?

OpenStudy (amistre64):

maybe .... hard to tell if i know how to add correctly or not these days

OpenStudy (anonymous):

or is it 2?

OpenStudy (amistre64):

2 is better, yes

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

i have one more question

OpenStudy (amistre64):

:) youre welcome

OpenStudy (anonymous):

Given the function f(x) = -3x3 + 9x2 - 2x + 3, what part of the function indicates that the left end starts at the top of the graph?

OpenStudy (anonymous):

1) The degree of the first term 2) The degree of the last term 3) The coefficient on the first term 4) The coefficient of the last term

OpenStudy (amistre64):

that question is not a good question. there is no 'start' of the graph unless something about it is restricted. and since different restrictions will produce different 'left side starts' the question is patently absurd.

OpenStudy (anonymous):

so theres no way to answer it?

OpenStudy (amistre64):

there is no way to answer it intelligently since the person that made the question evidently had no idea what they were asking for.

OpenStudy (amistre64):

spose i ask you: how do i get from here to there? what is missing about the situation that prevents you from giving a sensible response?

OpenStudy (anonymous):

the numbers

OpenStudy (amistre64):

well, more accurate start and ending values. the graph that your function produces is an infinite curve ... they have no start or end to them.

OpenStudy (anonymous):

if you had to make an educated guess what would you pick because i need to pick something

OpenStudy (amistre64):

2 things dictate the general shape of a polynomial: the coefficient of the highest degree, and the highest degree itself

OpenStudy (amistre64):

if i were to make a guess as to what in the world this question is talking about, id say that they assume that the domain and range of the graph is large enough, and that the 'parent' function is simply x^3. if the left side where to be positive as x approaches negative infinity, then the function would have to be adjusted as -x^3 so the sign of the leading coeff.

OpenStudy (anonymous):

okay thank you again

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks